OpenStudy (anonymous):

Katerina Lubov inherited some money and invested it at 6% simple interest.At the end of the year, the total amount of her original principal and the interest was $12,720.How much did she originally invest?

OpenStudy (nurali):

0.06x+x=12,720 1.06x=12,720 x=12,720 x=12,720/1.06 x=$12,000 was the original investment.

OpenStudy (nurali):

if correct please click best response.